OpenStudy (anonymous):

Estimate (a) how long it took King Kong to fall straight down from the top of the Empire State Building (380m high), and (b) his velocity just before "landing"?

OpenStudy (anonymous):

I know this is an easy problem... but the physics teach I have now is one of the worst teachers I have ever had. I NEED HELP!

OpenStudy (anonymous):

And if someone could explain this to me, that would be great! :)

OpenStudy (shane_b):

Part a: His initial velocity (Vi) is 0m/s and the only acceleration is due to gravity (9.8m/s^2). His time to fall can be calculated by using:\[d=V_it+\frac{1}{2}at^2\]\[380m=\frac{1}{2}(9.8m/s^2)t^2\]Solving for to you get t=8.81s. Part b: To find his final velocity (Vf) you can use this kinematic equation and plug in the time just just found above:\[V_f=V_i+at=(9.8m/s^2)(8.81s)=86.34m/s\]

OpenStudy (noelgreco):

Poor monkey!